Can I please get a walk through using a BA11 Plus financial calculator to this problem? I started, but got stuck.
30. A homeowner just obtained a $90,000 mortgage. The mortgage is for 30 years, with 360 monthly payment xed monthly interest rate of 9 12%. What percentage of the total payments made the first two years will go toward payment of interest? a. 89.30% b. 91.70% с. 92.59% d. 93.65% S72 pv:-90,000 e. 9476%
Answer #1

Answer:

Monthly Payment = $724.16
Monthly Interest Rate = 9/12% or 0.75%

Calculation of Loan outstanding after 2 years:

I/Y = 0.75%
n = 360 – 24 = 336
PMT = $724.16
Future Value = $0
PV = $88,712.52

Total Amount paid = 24 * $724.16
Total Amount paid = $17,379.84

Principal repaid = $90,000 - $88,712.52
Principal repaid = $1,287.48

Interest Paid = Total amount paid – Principal repaid
Interest Paid = $17,379.84 - $1,287.48
Interest Paid = $16,092.36

Percentage of Total payment made towards Interest = 16,092.36 / 17,379.84 * 100
Percentage of Total payment made towards Interest = 92.59%